# STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D Integration

### v1

@@ Abstract
Chiplet-based architectures and advanced packaging has emerged as transformative approaches in semiconductor design. While conventional physical design for 2.5D heterogeneous systems typically prioritizes wirelength reduction through tight chiplet packing, this strategy creates thermal bottlenecks and intensifies coefficient of thermal expansion (CTE) mismatches, compromising long-term reliability. Addressing these challenges requires holistic consideration of thermal performance, mechanical stress, and interconnect efficiency. We introduce STAMP-2.5D, the first automated floorplanning methodology that simultaneously optimizes these critical factors. Our approach employs finite element analysis to simulate temperature distributions and stress profiles across chiplet configurations while minimizing interconnect wirelength. Experimental results demonstrate that our thermal structural aware automated floorplanning approach reduces overall stress by 11% while maintaining excellent thermal performance with a negligible 0.5% temperature increase and simultaneously reducing total wirelength by 11% compared to temperature-only optimization. Additionally, we conduct an exploratory study on the effects of temperature gradients on structural integrity, providing crucial insights for reliability-conscious chiplet design. STAMP-2.5D establishes a robust platform for navigating critical trade-offs in advanced semiconductor packaging.
Index Terms:
Heterogeneous Integration, Automated Floor Planning, Chiplet Architecture, Thermal-Mechanical EvaluationI Introduction
@@
The continuous advancement of semiconductor technology has driven an ever-increasing demand for higher performance, greater functionality, and enhanced integration density in electronic systems. Chiplet-based architectures, particularly 2.5D integration, have emerged as promising solutions in the electronics industry, bridging the gap between chip and package sizes by mounting multiple dies on a single package substrate. This approach enhances capacity and performance by facilitating heterogeneous integration of chiplets with varied functionalities and material properties into unified systems [1]. Advanced packaging technologies, including Through-Silicon-Via (TSV) fabrication methods and multi-level assembly techniques, have further solidified the reliability and affordability of 2.5D packaging [2].
@@
Recent advancements in chiplet architectures and advanced packaging technologies have revolutionized system design, offering viable pathways to address the slowdown of Moore’s Law. Successful commercial implementations leverage chiplet architecture to integrate 7nm CPU dies with 12nm I/O dies for optimal performance and cost [3, 4]. Recent advances have further demonstrated scalability and efficiency in chiplet-based systems using heterogeneous 2.5D packaging techniques for workload-dependent configurations [5]. These innovations have catalyzed the development of diverse integration strategies, leveraging silicon interposers, bridge-chip technology, active interposers, and 3D interconnects [6, 1, 7]. Chiplet integration thus allows for scalable architectures, flexibility in node selection, and significant reuse of off-the-shelf intellectual properties (IPs)[8, 9]. However, substantial challenges persist, notably in inter-die communication, mixed-node integration, and thermal management[10, 11, 12].
@@
Thermal management has become increasingly critical in 2.5D chiplet systems, driven by higher integration densities and power densities in advanced semiconductor nodes. Extensive research efforts employing finite element analysis and computational fluid dynamics have studied thermal behaviors and factors influencing package warpage, die stress, and solder joint reliability [13, 2, 6]. Nevertheless, while thermal considerations have been extensively explored, mechanical stress implications, particularly arising from the mismatch in the Coefficient of Thermal Expansion (CTE) among heterogeneous dies and interposers, remain inadequately investigated. Thermomechanical stress poses substantial reliability risks in 2.5D packages, warranting integrated consideration alongside thermal and interconnect optimization [14].
To address these challenges, recent methodologies have started exploring chiplet placement strategies that incorporate thermal considerations. One such approach, TAP-2.5D, strategically inserts spacing between chiplets to jointly optimize thermal performance and wirelength [15]. However, existing methodologies predominantly focus on thermal and interconnect objectives, neglecting a critical component—the mechanical stresses due to thermal expansion mismatches—which significantly affect long-term reliability.
@@
In contrast, this paper presents STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D Integration, the first automated placement methodology that holistically integrates thermal performance, mechanical stress, and interconnect wirelength optimizations. Our proposed framework utilizes finite element analysis (FEA) simulations using ANSYS to concurrently evaluate thermal and mechanical stress profiles, enabling effective exploration of critical design trade-offs and ensuring robust reliability in advanced packaging.
@@ II Literature Review
Concurrently, automated placement methodologies have made significant strides in optimizing various design objectives, predominantly thermal performance, interconnect wirelength, and power efficiency. Prominent studies include those leveraging mathematical programming, heuristic-driven optimization, and machine-learning-based methods [20, 21, 15, 22, 23]. For instance, TAP-2.5D by Ma et al. (2023) explicitly introduced thermally-aware placement techniques, emphasizing insertion of spacing to reduce temperatures [15]. Chiou et al. (2023) and Deng et al. (2024) similarly utilized heuristic and reinforcement learning-based methods for thermal-aware floorplanning but did not explicitly incorporate mechanical stress as an objective or constraint in their methodologies [21, 23]. Although these studies represent significant advances in thermal and interconnect optimization, structural reliability and the critical implications of mechanical stress, especially arising from thermal gradients and interposer deformation, have been notably neglected. There is an critical need for automated placement methodology to jointly optimize chiplet configurations for thermal management, mechanical stress minimization, and interconnect wirelength efficiency.
@@ III-A Simulated Annealing
. This formulation accounts for the normalized change in length between iterations, ensuring that variations in dimension are factored into the cost function in a balanced manner relative to temperature and stress. The overall cost for a design is computed as a weighted sum of the normalized deviations in temperature, length, and stress. When normalization ranges are available, the cost functions for the current and candidate solutions are given by:
@@
If the normalization ranges are not applicable, a linear formulation is adopted instead. The difference in cost is defined as: . A positive (i.e., a decrease in cost) results in the immediate acceptance of the candidate solution. Conversely, if is negative, the candidate solution is accepted with a probability determined by the Boltzmann criterion: , where is the current annealing temperature. This probabilistic acceptance allows the algorithm to escape local minima by occasionally accepting solutions with higher costs, particularly during the early stages when is high. As the system cools, the likelihood of accepting suboptimal solutions diminishes, guiding the process toward convergence on a near-optimal configuration.
@@ III-B Von Mises Stress
where represents the von Mises stress, the terms , , and denote normal stresses in the respective coordinate directions, and , , and represent the shear stresses. Our approach employs this metric within ANSYS finite element analysis (FEA) to accurately capture structural stress profiles across heterogeneous chiplet arrangements.
@@ III-C Wirelength estimation
III-D Overview of Automated Floorplanning Algorithm
The proposed STAMP-2.5D methodology leverages a Simulated Annealing (SA) optimization framework to holistically optimize thermal management, structural reliability, and interconnect efficiency in chiplet-based systems. The algorithm initiates from an initial feasible chiplet placement—generated by adapting a widely-used automated floorplanning technique [24]—and iteratively explores improved configurations through stochastic perturbations.
@@ III-D Overview of Automated Floorplanning Algorithm
Fig. 1 provides a clear depiction of the iterative STAMP-2.5D floorplanning algorithm and highlights the integrated evaluation process.
@@ III-E Modeling
A refined mesh is applied in regions exhibiting steep thermal gradients and high stress concentrations—especially near interfaces and power-dense zones—to guarantee numerical precision. All interfaces (e.g., between chiplets and interposer, and between interposer and heat sink) are assumed to be perfectly bonded, thereby focusing the analysis on intrinsic material behavior while abstracting interfacial phenomena.
The simulation is conducted under steady-state conditions with a fixed ambient temperature of 23°C, and forced convective boundary conditions are applied on the heat sink surfaces to replicate realistic active cooling scenarios. By leveraging ANSYS’s multi-physics capabilities, our framework concurrently evaluates heat conduction and the resultant mechanical stresses. Specific experimental parameters, including localized power densities and variations in boundary conditions, are discussed in detail in the subsequent section.
@@ III-F Experimental Setup
Our experimental evaluation utilized three representative architectures: Micro150, Ascend910, and MultiGPU, which cover various chiplet configurations found in research and industry applications, as illustrated in Fig. 3. We implemented these architectures on a 2.5D package with a silicon interposer of dimensions 45mm × 45mm for ascend910 and 50mm x 50mm for the rest and thickness 100m. Each architecture consists of multiple chiplets with varying dimensions and functionalities.
For our thermal-mechanical-aware simulated annealing-based floor planning experiments, we maintained a consistent chiplet thickness of 150 across all chiplet types. This standardized thickness was chosen to represent typical manufacturing practices for high-performance chiplets. However, in our exploratory studies discussed in the results section, we varied the chiplet thickness to investigate the impact on thermal and mechanical performance. For thermal simulation, we applied non-uniform power density distributions across chiplets detailed in Fig. 3.
The boundary conditions for our thermal simulations included an ambient temperature of 23°C with forced convective cooling applied to the heat sink. We used architecture-specific heat transfer coefficients to model different cooling scenarios: 580 W/m²·K for Ascend910, 720 W/m²·K for Micro150, and 950 W/m²·K for MultiGPU [25]. These values reflect the different thermal requirements of each architecture, with higher-power systems requiring more effective cooling solutions. The bottom surface of the package substrate was set to a natural convection boundary condition with a heat transfer coefficient of 10 W/m²·K.
We performed all finite element analyzes using ANSYS Mechanical for static thermal structural analyzes, with a medium mesh consisting of approximately forty thousand elements, with particular refinement near interfaces and areas with high thermal gradients. The mesh convergence study ensured that our results were independent of mesh resolution.
Our optimization algorithm parameters included an initial annealing temperature of 1.0, a cooling rate of 0.95, and 50 iterations per temperature level, with a total of 4,500 iterations for each optimization run. These parameters were empirically determined to balance exploration and exploitation in the design space.
@@ IV Result and Analysis
We present experimental results from our thermal-mechanical aware floorplanning algorithm and analyze the critical relationship between temperature gradients and von Mises stresses in heterogeneous chiplet systems.
IV-A Thermal mechanical aware automated floor planning
We quantify temperature distributions, von Mises stress, and total interconnect wirelength through Ansys-based finite element analysis (FEA) simulations. Table II summarizes our results, clearly demonstrating the advantages of STAMP-2.5D compared to conventional thermal floor planning methodologies, while highlighting essential trade-offs among thermal management, structural reliability and connectivity efficiency.
Numerical results reveal compelling trade-offs. In the Micro150 architecture, our integrated approach notably reduces von Mises stress by 19.8%. This significant stress reduction is primarily driven by a symmetric chiplet placement strategy, which ensures uniform weight distribution across the interposer. Despite a modest temperature increase of only 2.97%, our method simultaneously achieves a wirelength reduction of 19.3%. The stress maps presented particularly in Fig. 4 clearly illustrate how symmetric placements lead to more balanced mechanical loads, contrasting sharply with the concentrated stress points typical of temperature-only optimization. This balanced weight distribution is critical in preventing localized solder joint failures and reducing interposer cracking during thermal cycling.
@@ IV-B Thermal Gradients and Structural Integrity Analysis
| Thermal Gradient Standard Deviation (Pa) | 466.11 | 495.40 | 410.59 | 932.96 | 1042.46 | 1050 | 708.47 | 711.46 | 653.01 |
| Hotspot Intensity Factor ((Tmax-Tavg)/Tavg) | 0.23 | 0.24 | 0.20 | 0.44 | 0.50 | 0.48 | 0.30 | 0.33 | 0.31 |
| Gradient-Stress Correlation Coefficient (r) | -0.25 | -0.22 | 0.02 | -0.25 | -0.15 | -0.15 | -0.22 | -0.24 | -0.24 |
| Stress Coefficient of Variance | 0.66 | 0.65 | 0.64 | 0.88 | 0.74 | 0.74 | 0.76 | 0.61 | 0.61 |
| Local Thermal-Mechanical Coupling Factor (Mean) (Pa.°C) | 59010.12 | 52912.96 | 33928.05 | 48198.98 | 39790.04 | 42687.99 | 34130.02 | 37143.14 | 40414.54 |
The peak stress that we observed in the previous subsection typically occurs on the interposer, highlighting the importance of analyzing how thermal gradients affect structural stability in our chiplet assemblies. Our analysis across three chiplet architectures reveals a fundamental insight: temperature gradients, rather than absolute temperatures, are the primary driver of mechanical stress in advanced chiplet assemblies. By incorporating mechanical stress considerations into our optimization framework, we achieve a sophisticated form of thermal-mechanical decoupling that enhances structural integrity without compromising performance.
The stress-aware optimization strategy consistently produces more uniform temperature gradient distributions across architectures, as illustrated in Fig. 7, which shows temperature gradient heat maps for all three optimization approaches. In Architecture 1 (Ascend910), this approach reduces the standard deviation of temperature gradients by nearly 12% and the mean gradient by 11.5%, creating more gradual thermal transitions that minimize stress concentrations. Architecture 3 (MultiGPU) shows similar improvements with an 8% reduction in gradient standard deviation. Remarkably, these improvements in gradient uniformity occur while maintaining nearly identical average temperatures around 65.6°C, confirming that the mechanical-thermal co-optimization methodology specifically targets harmful temperature discontinuities rather than merely reducing chip operating temperatures. The complete evaluation metrics for all three architectures are tabulated in Table III.
To quantify the degree of thermal-mechanical decoupling, we employ several specialized metrics. The Gradient-Stress Correlation Coefficient () measures the linear relationship between temperature gradients and mechanical stress, calculated as the Pearson correlation between and across all nodes in the finite element model. The Local Thermal-Mechanical Coupling Factor (LTMCF = ) quantifies how much stress is generated per unit of temperature gradient, with lower values indicating more effective decoupling of thermal effects from mechanical stress. The traditional Wirelength+Temperature optimization approach shows stronger negative Pearson correlations (around ) across all architectures, indicating that temperature gradients directly drive stress patterns. In contrast, the stress-informed floorplanning significantly weakens this correlation—shifting to near-zero (0.01) in Architecture 1 and reducing by over 40% in Architecture 2—demonstrating the ability to break the direct relationship between thermal gradients and mechanical stress.
Results from mechanical-aware optimization demonstrate promising stress reduction capabilities across architectures—with significant decreases of 19.8% in Micro150 and 7.8% in MultiGPU—despite maintaining or slightly increasing operating temperatures, indicating that temperature gradients likely play a more substantial role than absolute temperatures in thermal stress generation. Table III summarizes both thermal and stress metrics across all optimization strategies. These results clearly demonstrate that by explicitly considering mechanical stress during thermal optimization, chiplet designers can strategically manage temperature gradient distributions to minimize harmful stress concentrations, even in thermally challenging configurations. This mechanical-thermal co-optimization approach fundamentally transforms how thermal and mechanical behaviors interact in chiplet assemblies, offering a powerful strategy for enhancing long-term reliability in next-generation heterogeneous integration technologies.

### v2

@@ Abstract
Chiplet-based architectures and advanced packaging have emerged as transformative approaches in semiconductor design. While conventional physical design for 2.5D heterogeneous systems typically prioritizes wirelength reduction through tight chiplet packing, this strategy introduces thermal bottlenecks and intensifies coefficient of thermal expansion (CTE) mismatches, compromising long-term reliability. Addressing these challenges requires holistic consideration of thermal performance, mechanical stress, and interconnect efficiency. We introduce STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D integration, the first automated placement methodology that simultaneously optimizes these critical factors. Our approach employs finite element analysis to simulate temperature distributions and stress profiles across chiplet configurations while minimizing interconnect wirelength. Experimental results demonstrate that a thermal and structurally aware automated placement approach reduces overall stress by 11%, maintains excellent thermal performance with a negligible 0.5% temperature increase and simultaneously reduces total wirelength by 11% compared to temperature-only optimization. Additionally, we conduct an exploratory study on the effects of temperature gradients on structural integrity, providing crucial insights for reliability-conscious chiplet design.
I Introduction
@@
The continuous advancement of semiconductor technology has driven an ever-increasing demand for higher performance, greater functionality, and enhanced integration density in electronic systems. Chiplet-based architectures, particularly 2.5D integration, have emerged as promising solutions in the electronics industry, bridging the gap between chip and package sizes by mounting multiple dies on a single package substrate. This approach enhances capacity and performance by facilitating the heterogeneous integration of chiplets with varied functionalities and material properties into unified systems [1]. Advanced packaging technologies, including Through-Silicon-Via (TSV) fabrication methods and multi-level assembly techniques, have further solidified the reliability and affordability of 2.5D packaging [2].
@@
Recent advancements in chiplet architectures and advanced packaging technologies have revolutionized system design, offering viable pathways to address the slowdown of Moore’s Law. Commercial implementations leverage chiplet architecture to integrate 7nm CPU dies with 12nm I/O dies for optimal performance and cost [3, 4]. Further advances demonstrate scalability and efficiency in chiplet-based systems using heterogeneous 2.5D packaging techniques for workload-dependent configurations [5]. These innovations have catalyzed diverse integration strategies, including silicon interposers, bridge-chip technology, active interposers, and 3D interconnects [6, 1, 7]. Chiplet integration thus enables scalable architectures, flexibility in node selection, and reuse of off-the-shelf intellectual properties (IP) blocks[8, 9]. However, substantial challenges persist, notably in inter-die communication, mixed-node integration, and thermal management[10, 11, 12].
@@
Thermal management has become increasingly critical in 2.5D chiplet systemsdue to higher integration and power densities. One such approach, TAP-2.5D, strategically inserts spacing between chiplets to jointly optimize thermal performance and wirelength [13]. Extensive research efforts employing finite element analysis and computational fluid dynamics have studied thermal behaviors and factors influencing package warpage, die stress, and solder joint reliability [14, 2, 6]. Nevertheless, while thermal considerations have been extensively explored, mechanical stress implications, particularly arising from the mismatch in the Coefficient of Thermal Expansion (CTE) among heterogeneous dies and interposers, remain inadequately investigated. Beyond CTE-driven thermo-elastic stress, large 2.5-D assemblies are also subject to the self-weight of stacked chiplets, heat-spreaders, HBM modules and copper planes, which induces non-uniform bending and warpage across the interposer. In addition, residual stresses locked in during wafer thinning, die-attach curing and under-fill solidification can modify the absolute stress values as well. Because these mechanical loads differ in both magnitude and spatial distribution from the temperature field, thermomechanical stress poses substantial reliability risks in 2.5D packages, warranting integrated consideration alongside thermal and interconnect optimization [15].
Fig. 1 provides a simplified illustration of these two distinct stress drivers. In left panel, a uniform temperature field combined with asymmetric chiplet placement produces significant self-weight bending stress, whereas rigt panel shows how a symmetric layout with a high-power central die induces steep CTE-mismatch gradients and elevated thermo-elastic stress. These examples underscore the need to co-optimise temperature and mechanical stress.
@@
This paper presents STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D Integration, the first automated placement methodology that holistically integrates thermal performance, mechanical stress, and interconnect wirelength optimizations. Our proposed framework utilizes finite element analysis (FEA) simulations using ANSYS to concurrently evaluate thermal and mechanical stress profiles, enabling effective exploration of critical design trade-offs and ensuring robust reliability in advanced packaging. We also discuss how temperature gradients affect stress in the package and, finally, provide a fine-grained analysis of package temperature and stress distributions.
@@ II Literature Review
Concurrently, automated placement methodologies have made significant strides in optimizing various design objectives, predominantly thermal performance, interconnect wirelength, and power efficiency. Prominent studies include those leveraging mathematical programming, heuristic-driven optimization, and machine-learning-based methods [20, 21, 13, 22, 23]. For instance, TAP-2.5D by Ma et al. (2023) explicitly introduced thermally-aware placement techniques, emphasizing insertion of spacing to reduce temperatures [13]. Chiou et al. (2023) and Deng et al. (2024) similarly utilized heuristic and reinforcement learning-based methods for thermal-aware placement but did not explicitly incorporate mechanical stress as an objective or constraint in their methodologies [21, 23]. Although these studies represent significant advances in thermal and interconnect optimization, structural reliability and the critical implications of mechanical stress, especially arising from thermal gradients and interposer deformation, have been notably neglected. There is an critical need for automated placement methodology to jointly optimize chiplet configurations for thermal management, mechanical stress minimization, and interconnect wirelength efficiency.
@@ III-A Simulated Annealing
This formulation accounts for the normalized change in length between iterations, ensuring that variations in dimension are factored into the cost function in a balanced manner relative to temperature and stress. The overall cost for a design is computed as a weighted sum of the normalized deviations in temperature, length, and stress. When normalization ranges are available, the cost functions for the current and candidate solutions are given by:
@@
If the normalization ranges are not applicable, a linear formulation is adopted instead. The difference in cost is defined as: . A positive (i.e., a decrease in cost) results in the immediate acceptance of the candidate solution. Conversely, if is negative, the candidate solution is accepted with a probability determined by the Boltzmann criterion: .
The annealing temperature here is different from the package temperature. This probabilistic acceptance allows the algorithm to escape local minima by occasionally accepting solutions with higher costs, particularly during the early stages when is high. As the system cools, the likelihood of accepting suboptimal solutions diminishes, guiding the process toward convergence on a near-optimal configuration. The baselines for comparison are Wirelength + Temperature (WT) representing conventional thermal-aware placement and Wirelength + Stress (WS) optimizing mechanical reliability without thermal coupling. Both use identical SA parameters for fair comparison.
@@ III-B Von Mises Stress
where represents the von Mises stress, the terms , , and denote normal stresses in the respective coordinate directions, and , , and represent the shear stresses. Our approach employs this metric within ANSYS finite element analysis (FEA) to accurately capture structural stress profiles across heterogeneous chiplet arrangements. The FEA model applies both the steady-state temperature field and a gravity load (9.81 ms-2) to capture the self-weight induced shear stress by the components towards bending and warpage.
@@ III-C Wirelength estimation
III-D Overview of Automated placement Algorithm
The proposed STAMP-2.5D methodology leverages a Simulated Annealing (SA) optimization framework to holistically optimize thermal management, structural reliability, and interconnect efficiency in chiplet-based systems. The algorithm initiates from an initial feasible chiplet placement—generated by adapting a widely-used automated placement technique [24]—and iteratively explores improved configurations through stochastic perturbations.
@@ III-D Overview of Automated placement Algorithm
Fig. 2 provides a clear depiction of the iterative STAMP-2.5D placement algorithm and highlights the integrated evaluation process.
@@ III-E Modeling
A refined mesh is applied in regions exhibiting steep thermal gradients and high stress concentrations—especially near interfaces and power-dense zones—to guarantee numerical precision. The simulation is conducted under steady-state conditions with a fixed ambient temperature of 23°C, and forced convective boundary conditions are applied on the heat sink surfaces to replicate realistic active cooling scenarios. By leveraging ANSYS’s multi-physics capabilities, our framework concurrently evaluates heat conduction and the resultant mechanical stresses. Specific experimental parameters, including localized power densities and variations in boundary conditions, are discussed in detail in the subsequent section.
@@ III-F Experimental Setup
Our experimental evaluation utilized three representative architectures: Micro150, Ascend910, and MultiGPU, which cover various chiplet configurations found in research and industry applications, as illustrated in Fig. 4. We implemented these architectures on a 2.5D package with a silicon interposer of dimensions 45mm × 45mm for ascend910 and 50mm x 50mm for the rest and thickness 100m. Each architecture consists of multiple chiplets with varying dimensions and functionalities. For our thermal-mechanical-aware simulated annealing-based placement experiments, we maintained a consistent chiplet thickness of 150 across all chiplet types. This standardized thickness was chosen to represent typical manufacturing practices for high-performance chiplets. However, in our exploratory studies discussed in the results section, we varied the chiplet thickness to investigate the impact on thermal and mechanical performance. For thermal simulation, we applied non-uniform power density distributions across chiplets detailed in Fig. 4. The boundary conditions for our thermal simulations included an ambient temperature of 23°C with forced convective cooling applied to the heat sink. We used architecture-specific heat transfer coefficients to model different cooling scenarios: 580 W/m²·K for Ascend910, 720 W/m²·K for Micro150, and 950 W/m²·K for MultiGPU [25]. These values reflect the different thermal requirements of each architecture, with higher-power systems requiring more effective cooling solutions. The bottom surface of the package substrate was set to a natural convection boundary condition with a heat transfer coefficient of 10 W/m²·K.
All finite element analyses were run in ANSYS Mechanical for static thermomechanical conditions. We used a medium mesh of roughly forty thousand elements with local refinement near interfaces and strong thermal gradients, and we chose the coarsest mesh whose temperature predictions were negligible by additional refinement.
Our optimization algorithm parameters included an initial annealing temperature of 1.0, a cooling rate of 0.9, and 45 iterations per temperature level for Ascend910 and 50 for Micro150 and MultiGPU, with convergence at an annealing temperature of 0.01. This resulted in approximately 60.5 hours for Ascend910 and 67 hours for Micro150 and MultiGPU. With 5 independent runs for statistical validation, complete optimization requires 300-335 hours per architecture.
@@ IV Result and Analysis
We present experimental results from our thermal-mechanical aware placement algorithm and analyze the critical relationship between temperature gradients and von Mises stresses in heterogeneous chiplet systems. We compare our results against thermal-wirelength (TW) and thermal-stress (TS) co-optimisation only. The Thermal-Wirelength only optimisation is similar to [13] except that the geometry and consequently temperature prediction is performed using Ansys and not HotSpot.
IV-A Thermal mechanical aware automated placement
We quantify temperature distributions, von Mises stress, and total interconnect wirelength through Ansys-based finite element analysis (FEA) simulations. Table II summarizes our results, clearly demonstrating the advantages of STAMP-2.5D compared to conventional thermal placement methodologies, while highlighting essential trade-offs among thermal management, structural reliability and connectivity efficiency.
Numerical results reveal compelling trade-offs. In the Micro150 architecture, our integrated approach notably reduces von Mises stress by 19.8%. This significant stress reduction is primarily driven by a symmetric chiplet placement strategy, which ensures uniform weight distribution across the interposer. Despite a modest temperature increase of only 2.97%, our method simultaneously achieves a wirelength reduction of 19.3%. The stress maps presented particularly in Fig. 5 clearly illustrate how symmetric placements lead to more balanced mechanical loads, contrasting sharply with the concentrated stress points typical of temperature-only optimization. Because self-weight loading is included in the structural model, symmetric placement lowers bending moments on the interposer, thereby reducing peak even when peak rises slightly. This balanced load distribution is critical in preventing localized joint failures and reducing interposer cracking during thermal cycling.
@@ IV-B Thermal Gradients and Structural Integrity Analysis
| Mean Gradient | 871.56 | 968.90 | 774.59 | 1521.42 | 1546.38 | 1603.56 | 1457.60 | 1479.46 | 1322.73 |
| Thermal Gradient Standard Deviation | 487.60 | 503.18 | 444.88 | 956.72 | 1096.99 | 1202.41 | 879.24 | 819.58 | 735.29 |
| Max Gradient | 9286.08 | 6557.75 | 7384.52 | 16866.98 | 19009.43 | 18333.74 | 17955.89 | 15318.53 | 11880.72 |
| Temperature-Stress Correlation Coefficient (T-S) | -0.115 | -0.171 | -0.215 | -0.018 | -0.089 | -0.042 | -0.158 | -0.173 | -0.176 |
| Gradient-Stress Correlation Coefficient (G-S) | -0.252 | -0.236 | -0.246 | -0.256 | -0.173 | -0.173 | -0.217 | -0.220 | -0.222 |
The fundamental relationship shows thermal expansion induces stress proportional to temperature rise. At material interfaces, this becomes , where CTE mismatches amplify stress. However, stress generation in 2.5D chiplet systems involves multiple loading mechanisms beyond thermal expansion alone.
Our analysis confirms that temperature gradients, rather than absolute temperatures, are the primary driver of thermally-induced mechanical stress. As shown in Table III, the gradient-stress correlation coefficients (G-S Corr: to ) are consistently stronger than temperature-stress correlations (T-S Corr: to ) across all architectures and optimization approaches. This validates that spatial temperature variations create the thermal expansion mismatches that generate mechanical stress.
Including temperature and stress in the objective leads the SA framework to reduce thermal gradients and mechanical loading asymmetries. Consequently, placements tend to show reduced co-location of temperature and stress hotspots, although full decoupling is not guaranteed. The stress-aware optimization strategy consistently produces more uniform temperature gradient distributions across architectures, as illustrated in Fig. 8, which shows temperature gradient heat maps for all optimization approaches. In Ascend910, our integrated WST approach significantly reduces gradient standard deviation ( vs , reduction) and maximum gradient ( vs , reduction) as detailed in Table III. MultiGPU demonstrates similar improvements with gradient standard deviation reducing from to ( reduction) and maximum gradient decreasing from to ( reduction). These gradient reductions, clearly visible in the gradient maps of Fig. 8, directly correlate with the observed stress improvements while maintaining effective thermal management.
The Micro150 architecture provides particularly compelling evidence for multi-physics stress generation beyond thermal effects. Despite thermal-only optimization (TW) achieving relatively controlled thermal gradients (Std Grad: , Table III), it produces high absolute stress values due to asymmetric chiplet placement that creates unbalanced self-weight loading across the interposer. This asymmetric mass distribution generates bending moments independent of thermal effects, demonstrating why comprehensive thermal-mechanical co-optimization is essential for architectures with non-uniform chiplet arrangements.
IV-C Impact of Intra-Chiplet Power Density on Thermal-Mechanical Reliability
To assess the effect of finer-grained power variation on placement quality, we take the final STAMP-2.5D placements generated under each objective set—wirelength+temperature (WT), wirelength+stress (WS) and the full multi-objective (WST) for the Vitruvian chiplet in the Ascend910 system and recompute their steady-state temperature and von Mises stress fields for a spatially varying heat-flux profile derived from the block-level power map of the Vitruvian compute die (Figure 1)[28]. By holding placement constant and only changing the power input, this study isolates how intra-die hotspots influence package‐level thermal and mechanical behavior without retraining or retuning the optimiser.
| Configuration | Uniform | Non-uniform | ||
| Temp (°C) | Stress (MPa) | Temp (°C) | Stress (MPa) | |
| WT | 90.99 | 256.98 | 91.17 | 258.82 |
| WS | 91.27 | 249.91 | 91.48 | 251.45 |
| WST | 87.63 | 242.65 | 89.08 | 243.29 |
Introducing the non-uniform hotspot map raises absolute peak temperatures and stresses only marginally (Table IV). Crucially, the relative ranking of the three placement modes remains identical under both power assumptions: WST produces the lowest peak temperature and , followed by WT, then WS. Figure 10 shows the temperature and stress contours for the WST placement under the hotspot model; WT and WS follow the similar trends. Although these findings confirm that STAMP-2.5D’s optimal placements are robust to realistic intra-chiplet power variation, implying that a simple average-power model suffices for early‐stage thermal–mechanical placement, a more granular examination of their transient behaviors remains as planned future work to further validate and refine the methodology.